VIBRATIONAL AND GEOMETRIC STRUCTURES OF La_{3}C_{2}O AND La_{3}C_{2}O^{+} FROM MASSE-ANALYZED THRESHOLD IONIZATION

\underline{ROUDJANE MOURAD}, LU WU and D. S. YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

La_{3}C_{2}O is produced for the first time by laser vaporization in a pulsed cluster source and identified by photoionization time-of-flight mass spectrometry. Vibrationally-resolved ion spectra are obtained with mass-analyzed threshold ionization (MATI) spectroscopy. The adiabatic ionization energy of La_{3}C_{2}O is measured to be 30891(5) cm^{−1}. The spectra display several short vibrational progressions, and these progressions are associated mainly with La-La, La-C and La_{3}C_{2}O stretching excitations. The electron-spin multiplicities and molecular symmetries of La_{3}C_{2}O and La_{3}C_{2}O^{+} are determined by combining the experimental measurements with ab initio calculations at MP2 level. Preliminary data analysis shows that the ^{1}A_{1} ← ^{2}A_{1} transition is responsible for the observed MATI spectra. The cluster has C_{2v} symmetry with La_{3}C_{2}O in a bi-pyramid structure and oxygen being attached to the La_{3} plane.